OpenStudy (anonymous):

solve the following system of linear equations. please verify your solution. 4x + 3y = 13 & 5x + y - 8 = 0

hero (hero):

4x + 3y = 13 5x + y = 8 y = (13 - 4x)/3 y = 8 - 5x y = y 13 - 4x = 3(8 -5x) 13 - 4x = 24 - 15x 11x = 11 x = 11/11 x = 1 y = 3